OpenStudy (scorcher219396):

Differentiate using chain rule: y=cos(a^3+x^3)

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{midnightblue}{f(x)=cos(a^3+x^3)\\ f'(x)=cos'(a^3+x^3)\times(a^3+x^3)'}\)

OpenStudy (xapproachesinfinity):

Now what is the derivative of cos? derivative of (x^3)' should be easy to you

OpenStudy (scorcher219396):

Right, it's -sin and 3x^2 But what happened to the a's?

OpenStudy (xapproachesinfinity):

@Disturb no answer pls, let him do it

OpenStudy (scorcher219396):

(or her) :)

OpenStudy (xapproachesinfinity):

a^3 is constant term

OpenStudy (scorcher219396):

Oh ok

OpenStudy (scorcher219396):

That makes sense

OpenStudy (xapproachesinfinity):

well i forgot to put her too may bad^_^

OpenStudy (xapproachesinfinity):

well i guess you figured out the answer now?

OpenStudy (scorcher219396):

Yep! Got it! thanks!

OpenStudy (xapproachesinfinity):

Anytime!

OpenStudy (anonymous):

not

OpenStudy (anonymous):

me got suspended